OpenStudy (anonymous):

John deposited $10,000 today in a bank where the interest rate is 4%. (a) How much will he have after 7 years if the interest is compounded quarterly? (b) How much will he have after 7 years if the interest is compounded continuously? (c) How long does it take to triple his money if the interest is compounded continuously? ****please explain how you got your answers***

OpenStudy (r.hamza17):

Do you have the concept of compound interest?

OpenStudy (anonymous):

no that's whats confusing me

OpenStudy (r.hamza17):

Okay. To find compound interest formula used is A=P(a\1+r)^n I=A-P

OpenStudy (anonymous):

ok

OpenStudy (r.hamza17):

But when it is given compounded quarterly than you have to multiply 4 to n and divide 4 from r as their are 4 quarters in one year

OpenStudy (anonymous):

im still confused could you write out the formula for me

OpenStudy (r.hamza17):

The formula is the same but you can derive it as follows for compounded quarterly like this; A=P(1+r/2)n/2

OpenStudy (anonymous):

but how do I plug it in?

OpenStudy (r.hamza17):

You have got the values in your question n=7 years P=$10000 r=4%=0.04 put these values in the above formula

OpenStudy (anonymous):

ok what about for parts b and c?

OpenStudy (anonymous):

does the formula change?

OpenStudy (r.hamza17):

In part b you should use the simple formula\[A=Pe ^{rt}\]

OpenStudy (r.hamza17):

And the same one in c also

OpenStudy (anonymous):

what gets plugged in for t? time?

OpenStudy (r.hamza17):

In compound interest t and n have the same meaning time. Here t/n=7

OpenStudy (anonymous):

oh ok thankyou